OpenStudy (anonymous):

create your own radical expression problem with the answer and three sentences about radical expressions MEDALS TO ALL

OpenStudy (anonymous):

will assume you meant √ ( 20 x^6 y^3) = 2 y | x^3 |√ ( 5y ) note : if x = -2 then [ -2 ]^6 = 64 , but √64 = 8 = 2 ^3 , not (-2)^3...thus |...| are used but y^3 > 0 means y > 0 . so |...| not needed